In the following fiNine students are seated evenly around a circular table. They are
marked on the circumference of the circle as A, B, C, D, E, F, G, H, I.
If the center of the circle is situated at point O, what is m<OAD
1
Expert's answer
2020-11-22T13:25:22-0500

"\displaystyle\textsf{Since, they are seated evenly}\\
\textsf{around a circular table,}\\


\textsf{Distance between}\, A, B, C, D,...,I \,\, \textsf{is}\\
\frac{360\degree}{9} = 40\degree\\


m<AOD = 40\degree \times 3 = 120\degree\\


m<AOD + m<OAD + m<ODA = 180\degree\\

m<ODA = m<OAD\\


120\degree + 2m<OAD = 180\degree\\

2m<OAD = 60\degree\\

m<OAD = 30\degree"